(a) If we roll two fair dice, what is the probability that a total of nine shows?
(b) if we toss four fair coins, what is the probability that we get exactly two heads?

Answer by checkley71(8403)

TOTAL OUTCOMES=36 (6*6) WAYS TO GET NINE ARE 3&6 4&5 5&4 6&3 OR 4 WAYS TO GET NINE TOTAL THEREFOIRE 4/36 = 1/8 OR 12.5% OF GETTING A 9
TOTAL OUTCOMES=16 AND 6 WILL HAVE 2 HEADS THUS HHTT, HTHT, HTTH, THHT, TTHH, THTH OR 6/16 OR 37.5%